OpenStudy (anonymous):

Solve systematic equation using both the substitution method and the addition method. x=3y-7 2x-6y=-14

OpenStudy (anonymous):

Could you kindly show me step by step?

OpenStudy (anonymous):

I tried plugging what equals x into the second equation but I am stuck.

OpenStudy (anonymous):

2(3y-7)-6y=-14 6y-14-6y=-14 -14 = -14. Get it?

OpenStudy (anonymous):

That's the substitution method?

OpenStudy (anonymous):

Yes, I got that. How do I do the addition method?

OpenStudy (anonymous):

isn't it 0=0 when you substitute x into the equation.

OpenStudy (anonymous):

Yes, you are right. It is 0 = 0 for the substitution method. For the addition method: x = 3y -7 2x - 6y = -14 ------------ x - 3y = -7 2x - 6y = -14 ------------ -2x + 6y = 14 2x - 6y = -14 ------------ Then add, you get 0 = 0 as well.

OpenStudy (anonymous):

Like I showed you earlier, -14 = -14 which tells us that both equations are the same regardless.

OpenStudy (anonymous):

Oh okay, I understand. Thank you :)